# Formulating or Fixating: Effects of Examples on Problem Solving Vary as a Function of Example Presentation Interface Design

### 19 Jan 2024 v1

@@ 1. Introduction
Examples — descriptions or representations of possible solutions (or parts thereof) for the same or related problems (Sio et al., 2015; Crilly and Cardoso, 2017; Herring et al., 2009; Smith et al., 1993; Brandt et al., 2010a) — are an integral part of the creative problem solving process. Examples can take many forms, such as previous physical prototypes brought to a brainstorm (Hargadon and Sutton, 1997), search results for patents for related problems (Fu et al., 2013), spoken ideas from collaborators (Nijstad and Stroebe, 2006), UI designs in an accessible gallery (Kumar et al., 2011)), or references from memory to earth animals or science fiction creatures when inventing new fictional alien creatures (Ward, 1994)). Importantly, examples can substantially shape what ideas come to mind (Ward, 1994; Smith et al., 1993). This ”structuring of imagination” (Ward, 1994) is sometimes helpful “inspiration” that leads to more creative ideas (Eckert and Stacey, 1998; Dahl and Moreau, 2002; Hargadon and Sutton, 1997; Siangliulue et al., 2015a). But examples can also have harmful “fixation” effects that constrain novelty and innovation (Jansson and Smith, 1991; Linsey et al., 2010; Ward, 1994; Smith et al., 1993) (for recent reviews, see (Sio et al., 2015) and (Crilly and Cardoso, 2017)). Importantly, examples can influence problem solving without conscious effort or recognition (Marsh et al., 1999; Marsh and Bower, 1993; Linsey et al., 2010), and persist in spite of creators’ explicit intentions not to be influenced by them (Smith et al., 1993; Vasconcelos et al., 2018; Jansson and Smith, 1991). Perhaps in recognition of these facts, effective creators take an active role in finding, structuring and interacting with examples (Eckert and Stacey, 1998; Herring et al., 2009; Sharmin et al., 2009) using a variety of analog and digitally mediated systems and practices, such as search engines (Herring et al., 2009; Zhang et al., 2020; Palani et al., 2021b; Palani et al., 2021a), design workbooks and commonplace notebooks (Gaver, 2011), online whiteboards (Wallace et al., 2020), mood boards (Lucero, 2012), and wider interactions with their community of practice, such as trade publications and conventions (Eckert and Stacey, 1998; Hargadon and Sutton, 1997).
@@ 2.2. Theoretical insights into human-example interaction
A number of detailed in-situ studies of creators have documented a range of strategies for working with examples, ranging from simpler, more source-driven strategies like direct source adaptation (Eckert and Stacey, 2003), to more complex and reflective strategies associated with more radical transformation of examples, such as source analysis and schema-driven source selection (Eckert and Stacey, 2003), analogical reasoning (Gentner and Markman, 1997; Holyoak and Thagard, 1996; Gick and Holyoak, 1983; Ball et al., 2004), and generating novel emergent features that can connect disparate attributes across examples (Wilkenfeld and Ward, 2001). These “processing strategies” can be described by a variety of theoretical frames from the psychological literature on creativity. We believe this theoretical level of description could facilitate our goal of synthesizing mappings between interface characteristics and effects of examples on creative problem solving outcomes. Some notable examples include basic memory mechanisms such as priming (McNamara, 1992) and spreading activation (Collins and Loftus, 1975; Raaijmakers and Shiffrin, 1981), and higher level cognitive processes such as conceptual abstraction and analogical transfer (Gick and Holyoak, 1983), conceptual combination (Wilkenfeld and Ward, 2001), and problem reformulation based on examples (Helms et al., 2008; Dorst and Cross, 2001). Of particular interest in our study is a contrast between priming and spreading activation mechanisms on the one hand, which is associated with lower-level conceptual influences, and problem (re)formulation, which is associated with more complex, higher-order processing of examples. In this study, we extend this literature by exploring how two specific mechanisms of processing examples (stimulation, and (re)formulation) might be helped or hindered by different example interaction interfaces. To set the context for our results, we briefly review the literature on each mechanism here.
@@ 2.2.2. Using examples to (re)formulate problems
Past research has documented how people can use examples to construct, refine, and even reformulate their understanding of the creative problem they are trying to solve (Helms et al., 2008; Okada et al., 2009; Herring et al., 2009; Sharmin et al., 2009; Maier, 1931; Kaplan and Simon, 1990), through processes such as intentional free-form curation of examples (Lupfer et al., 2016) or on mood boards (Lucero, 2012). For instance, Okada et al documented how two artists used individual artworks to shape not just their ideas, but used a process of “analogical modification” to search for and modify higher-order concepts, and their creative vision over the course of years (Okada et al., 2009). This process of example-influenced problem formulation is related to computational and neurobiological models of cognitive search (Hills et al., 2015)which study the range of strategies that intelligent agents can use to structure their search processes: here, there is an important distinction between “model-free” search, where external feedback from the world on the agents’ actions guide search in a simpler, more local, stimulus-response manner, and “model-based” search, where the agent constructs a model or representation of the task and environment (in the case of creative problem-solving, this would be the problem and/or design space (Newell and Simon, 1972; Goel and Pirolli, 1992; Dorst and Cross, 2001)) partly on the basis of reflection its own actions and possibly observation of others’ actions, and uses that model to decide where and when to explore in the task environment vs. continue to sample locally. Additionally, insight problem solving research has documented how people not only construct models, but also substantially revise them in radical ways, to solve difficult creative problems (Knoblich et al., 1999; Kaplan and Simon, 1990); this process is a key source of difficulty for creative problems, where one’s initial problem formulation (e.g., key constraints or requirements), may be unhelpful (Knoblich et al., 1999). In this paper, we discuss this set of mechanisms under the label “(re)formulation”, to capture the intuition of creators leveraging examples to (re)formulate their understanding of the design space.
@@ 3.1.2. Examples
To prepare sets of initial examples with High Diversity (HD) or Low Diversity (LD), we randomly generated 10,000 sets of 10 examples each (recall that each example is a ”drilling location” point in the 100x100 Wildcat Wells search environment, with a corresponding score) for each of our 10 search environments, and ranked the diversity of each example set with a close variant of the Determinantal Point Process (DPP) approach (Kulesza and Taskar, 2012) described in Algorithm 2 in Appendix B; intuitively, this approach measured the “hyper-volume” spanned by a selected set of points, such that larger volumes corresponded to higher levels of diversity, since these points spanned a larger set of the space of possible moves (Kulesza and Taskar, 2012). We then randomly picked three HD examples sets that were greater than the 99th percentile of the distribution of diversity across the example sets, and three LD examples sets with diversity lower than 1st percentile of the diversity distribution. To ensure that examples would not directly reveal the location of the peak, or provide a high enough score that participants might simply stop after seeing the example instead of searching, we discarded example sets that had any point with a score over 80 (scores in the search environment ranged from 0 to 100), and resampled example sets as necessary — subject to the same low/high diversity sampling criteria) to construct our final example sets for each search environment. Figure 1 (B) shows an LD and HD example set used in our experiment.

### 23 Jan 2024 v2

@@ 1. Introduction
Examples — descriptions or representations of possible solutions (or parts thereof) for the same or related problems (Sio et al., 2015; Crilly and Cardoso, 2017; Herring et al., 2009; Smith et al., 1993; Brandt et al., 2010a) — are an integral part of the creative problem solving process. Examples can take many forms, such as previous physical prototypes brought to a brainstorm (Hargadon and Sutton, 1997), search results for patents for related problems (Fu et al., 2013), spoken ideas from collaborators (Nijstad and Stroebe, 2006), UI designs in an accessible gallery (Kumar et al., 2011)), or references from memory to earth animals or science fiction creatures when inventing new fictional alien creatures (Ward, 1994)). Importantly, examples can substantially shape what ideas come to mind (Ward, 1994; Smith et al., 1993). This ”structuring of imagination” (Ward, 1994) is sometimes helpful “inspiration” that leads to more creative ideas (Eckert and Stacey, 1998; Dahl and Moreau, 2002; Hargadon and Sutton, 1997; Siangliulue et al., 2015a). But examples can also have harmful “fixation” effects that constrain novelty and innovation (Jansson and Smith, 1991; Linsey et al., 2010; Ward, 1994; Smith et al., 1993) (for recent reviews, see (Sio et al., 2015) and (Crilly and Cardoso, 2017)). Importantly, examples can influence problem solving without conscious effort or recognition (Marsh et al., 1999; Marsh and Bower, 1993; Linsey et al., 2010), and persist in spite of creators’ explicit intentions not to be influenced by them (Smith et al., 1993; Vasconcelos et al., 2018; Jansson and Smith, 1991). Perhaps in recognition of these facts, effective creators take an active role in finding, structuring and interacting with examples (Eckert and Stacey, 1998; Herring et al., 2009; Sharmin et al., 2009) using a variety of analog and digitally mediated systems and practices, such as search engines (Herring et al., 2009; Zhang et al., 2020; Palani et al., 2021b; Palani et al., 2021a), design workbooks and commonplace notebooks (Gaver, 2011), online whiteboards (Wallace et al., 2020; MacNeil et al., 2023), mood boards (Lucero, 2012), and wider interactions with their community of practice, such as trade publications and conventions (Eckert and Stacey, 1998; Hargadon and Sutton, 1997).
@@ 2.2. Theoretical insights into human-example interaction
A number of detailed in-situ studies of creators have documented a range of strategies for working with examples, ranging from simpler, more source-driven strategies like direct source adaptation (Eckert and Stacey, 2003), to more complex and reflective strategies associated with more radical transformation of examples, such as source analysis and schema-driven source selection (Eckert and Stacey, 2003), analogical reasoning (Gentner and Markman, 1997; Holyoak and Thagard, 1996; Gick and Holyoak, 1983; Ball et al., 2004), and generating novel emergent features that can connect disparate attributes across examples (Wilkenfeld and Ward, 2001). These “processing strategies” can be described by a variety of theoretical frames from the psychological literature on creativity. We believe this theoretical level of description could facilitate our goal of synthesizing mappings between interface characteristics and effects of examples on creative problem solving outcomes. Some notable examples include basic memory mechanisms such as priming (McNamara, 1992) and spreading activation (Collins and Loftus, 1975; Raaijmakers and Shiffrin, 1981), and higher level cognitive processes such as conceptual abstraction and analogical transfer (Gick and Holyoak, 1983; Ding et al., 2023), conceptual combination (Wilkenfeld and Ward, 2001), and problem reformulation based on examples (Helms et al., 2008; Dorst and Cross, 2001; MacNeil et al., 2021). Of particular interest in our study is a contrast between priming and spreading activation mechanisms on the one hand, which is associated with lower-level conceptual influences, and problem (re)formulation, which is associated with more complex, higher-order processing of examples. In this study, we extend this literature by exploring how two specific mechanisms of processing examples (stimulation, and (re)formulation) might be helped or hindered by different example interaction interfaces. To set the context for our results, we briefly review the literature on each mechanism here.
@@ 2.2.2. Using examples to (re)formulate problems
Past research has documented how people can use examples to construct, refine, and even reformulate their understanding of the creative problem they are trying to solve (Helms et al., 2008; Okada et al., 2009; Herring et al., 2009; Sharmin et al., 2009; Maier, 1931; Kaplan and Simon, 1990), through processes such as intentional free-form curation of examples (Lupfer et al., 2016) or on mood boards (Lucero, 2012). For instance, Okada et al. documented how two artists used individual artworks to shape not just their ideas, but used a process of “analogical modification” to search for and modify higher-order concepts, and their creative vision over the course of years (Okada et al., 2009). This process of example-influenced problem formulation is related to computational and neurobiological models of cognitive search (Hills et al., 2015) which study the range of strategies that intelligent agents can use to structure their search processes: here, there is an important distinction between “model-free” search, where external feedback from the world on the agents’ actions guide search in a simpler, more local, stimulus-response manner, and “model-based” search, where the agent constructs a model or representation of the task and environment (in the case of creative problem-solving, this would be the problem and/or design space (Newell and Simon, 1972; Goel and Pirolli, 1992; Dorst and Cross, 2001)) partly on the basis of reflection its own actions and possibly observation of others’ actions, and uses that model to decide where and when to explore in the task environment vs. continue to sample locally. Additionally, insight problem solving research has documented how people not only construct models, but also substantially revise them in radical ways, to solve difficult creative problems (Knoblich et al., 1999; Kaplan and Simon, 1990); this process is a key source of difficulty for creative problems, where one’s initial problem formulation (e.g., key constraints or requirements), may be unhelpful (Knoblich et al., 1999). In this paper, we discuss this set of mechanisms under the label “(re)formulation”, to capture the intuition of creators leveraging examples to (re)formulate their understanding of the design space.
@@ 3.1.2. Examples
To prepare sets of initial examples with High Diversity (HD) or Low Diversity (LD), we randomly generated 10,000 sets of 10 examples each (recall that each example is a ”drilling location” point in the 100x100 Wildcat Wells search environment, with a corresponding score) for each of our 10 search environments, and ranked the diversity of each example set with a close variant of the Determinantal Point Process (DPP) approach (Kulesza and Taskar, 2012) described in Algorithm 2 in Appendix B proposed by Eesh et al. (Kamrah et al., 2023); intuitively, this approach measured the “hyper-volume” spanned by a selected set of points, such that larger volumes corresponded to higher levels of diversity, since these points spanned a larger set of the space of possible moves (Kulesza and Taskar, 2012). We then randomly picked three HD examples sets that were greater than the 99th percentile of the distribution of diversity across the example sets, and three LD examples sets with diversity lower than 1st percentile of the diversity distribution. To ensure that examples would not directly reveal the location of the peak, or provide a high enough score that participants might simply stop after seeing the example instead of searching, we discarded example sets that had any point with a score over 80 (scores in the search environment ranged from 0 to 100), and resampled example sets as necessary — subject to the same low/high diversity sampling criteria) to construct our final example sets for each search environment. Figure 1 (B) shows an LD and HD example set used in our experiment.